OpenStudy (anonymous):

Divide and simplify Having some problems with this one. 2w-18/121 ÷ w-9/55w

OpenStudy (anonymous):

I know i turn it into 2w-18/121 * 55w/w-9 but after that I feellike im missing a step

OpenStudy (hoblos):

2w-18 =2(w-9)  =10w/11